OpenStudy (anonymous):

Express each product in the simplest form. 3wx / 6x times 3wx / 9w a. wx / 6 b. wx / 3 c. 2 / 3 WX d. wx Help me please. :)

OpenStudy (imstuck):

\[\frac{ 3wx }{ 6x } * \frac{ 3wx }{ 9w }\]You can cross cancel here. Do you know how?

OpenStudy (somy):

\[ \frac{ 3wx }{ 6x }\times \frac{ 3wx }{ 9w }\] simplest form for 3wx/6x would be w/2 x cancel out and 3 and 6 are divided by 3 for 3wx/9w simplest form will be x/3 w is cancelled out and 3 and 9 are divided by 3 so u are left with \[\frac{ w }{ 2 }\times \frac{ x }{3 }\] so it'll be wx/6

OpenStudy (anonymous):

Thank you both!

OpenStudy (anonymous):

I will study your responses.

OpenStudy (somy):

:)

OpenStudy (imstuck):

If you want to do each fraction at a time, then cross cancel between the 3 and the 6 and the x's to get a result of \[\frac{ w }{ 2 }\]Then work with the next fraction to eliminate between the 3 and the 9, and the w's to get a result of\[\frac{ x }{ 3}\]

OpenStudy (imstuck):

Now you have, simply,\[\frac{ w }{ 2 } * \frac{ x }{ 3 }= \frac{ wx }{ 6 }\]

OpenStudy (anonymous):

Thank you I am stuck I wish I could give a second medal :C

OpenStudy (somy):

@Soccergirl52 i gave him/her :)